OpenStudy (anonymous):

Hi, Can any body help me? I have xy coordinates of a line (like 10 points), how could I calculate the length of the line

OpenStudy (amistre64):

add up the distances from point to point for a good approximation

OpenStudy (amistre64):

then divvy up you interval into an infinite number of partitions and measure each point

OpenStudy (anonymous):

Divvy means?

OpenStudy (amistre64):

by line, you mean a straight line? or just some curve drawn in pencil?

OpenStudy (amistre64):

divide

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

curved line

OpenStudy (amistre64):

\[\int_{a}^{b}\sqrt{1+[f'(x)]^2}dx\]

OpenStudy (anonymous):

what is this? can you explain me?

OpenStudy (amistre64):

if you just have the 10 points; then the distance from point to point added up should be fine

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

thats calculus; its the divving up of the interval from a to b an infinite number of times and adding up the points

OpenStudy (amistre64):



OpenStudy (anonymous):

So I can use that formula to find the length.

OpenStudy (amistre64):

you can, but its not what they want you to do :) the distance from point to point is thge pythag thrm

OpenStudy (amistre64):

\[x^2+y^2 = s^2\] \[s = \sqrt{x^2+y^2}\]

OpenStudy (amistre64):

and of course x and y here stand for the distance between the points

OpenStudy (amistre64):

you got 2 points I can give an example on?

OpenStudy (anonymous):

so, x is the first point (x1,y1) and y is the second point coordinates (x2,y2) X = (x1-x2), Y=(y1-y2) is this correct?

OpenStudy (amistre64):

yeah, thats good

OpenStudy (anonymous):

then I have to add all s1,s2,s3

OpenStudy (amistre64):

yep

OpenStudy (amistre64):

otherwise it be kinda hard to determine what they sum up to be :)

OpenStudy (anonymous):

I have another question. Can you help me?

OpenStudy (amistre64):

wish i could, but I gotta run. good luck :)

OpenStudy (anonymous):

ok thanks